OpenStudy (anonymous):

brianna plotted the two points (20, 75) and (45, 150) on a graph .what is the x-coordinate of the x-intercept of the line that contains these two points ?

OpenStudy (jdoe0001):

first off, find the equation of the line \(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &(20\quad ,&75)\quad &(45\quad ,&150) \end{array} \\\quad \\ slope = m= \cfrac{rise}{run} \implies \cfrac{y_2-y_1}{x_2-x_1} \\ \quad \\ y-y_1=m(x-x_1)\qquad \textit{plug in your values and solve for "y"}\) once you get the equation of the line, to find the x-intercept, just set y = 0, and solve for "x"